Define ~ on R by ‘ a ~ b iff a − b∈Z ’. Check whether or not ~ is an equivalence relation on R. If it is, find [√5 ]Else, give another equivalence relation on R
1
Expert's answer
2020-03-13T09:48:42-0400

Reflexive: if "a \in R" then "a-a=0\in Z" . Hence "\sim" is reflexive.

Symmetric: if "a,b\in R \space and \space a-b\in Z" then "b-a" also belongs to Z

Hence "\sim" is symmetric.

Transitive: If "a,b,c\in R \space and \space a-b\in Z,b-c\in Z \space then \space 
a-b=n" for some "n\in Z \space and \space b-c=m" for some "m\in Z\space then \space a=n+b \space and \space c=b-m
\implies a-c=(n+b)-(b-m)=n-m\in Z ."

Therefore "\sim" is transitive.

Hence "\sim" is an equivalence relation on R.


["\sqrt{5}"]"=" { "x\in R :x\sim \sqrt{5}" }

"=" { "x\in R:x-\sqrt{5} \in Z \space i,e \space x-\sqrt{5}=n" for some "n\in Z" }

"=" { "x\in R: x=n+\sqrt{5}" where "n\in Z" }